# Gravity as a topological gauge theory

### v3

@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}=\overbrace{\varpi_{1}\oplus\varpi_{0}}^{\varpi_{\mathfrak{h}}}\oplus\varpi_{-1}=$ $\varpi_{1}^{a}\begin{pmatrix}0&\bar{e}_{a}&0\\ 0&0&e_{a}\\ 0&0&0\end{pmatrix}+\begin{pmatrix}\lambda&0&0\\ 0&A&0\\ 0&0&-\lambda\end{pmatrix}+\varpi_{-1}^{a}\begin{pmatrix}0&0&0\\ e_{a}&0&0\\ 0&\bar{e}_{a}&0\end{pmatrix}$
@@
where $A=\dfrac{1}{2}A^{ab}J_{ab}=\dfrac{1}{2}A^{ab}_{\mu}dx^{\mu}\otimes J_{ab}$ corresponds to the spin connection and $\beta=\beta^{a}e_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes e_{a}$ corresponds to the tetrad or soldering form used to define the metric of the tetrad formalism $\tilde{g}=\eta(\beta,\beta)$. $\alpha=\alpha^{a}e_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes e_{a}$ can be interpreted as a secondary tetrad. Finally $\lambda=\lambda_{\mu}dx^{\mu}$ is the $\mathbb{R}$-valued (dilations) part of the connection while $\ell$ and $\ell^{\prime}$ are non null real numbers. Also : $\bar{\alpha}=\alpha^{a}\bar{e}_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$ and $\bar{\beta}=\beta^{a}\bar{e}_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$.
@@
$\bar{\Omega}$ $=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{p}}=\begin{pmatrix}f&\dfrac{1}{\ell^{\prime}}\bar{\Pi}&0\\ \dfrac{1}{\ell}\Theta&F&\dfrac{1}{\ell^{\prime}}\Pi\\ 0&\dfrac{1}{\ell}\bar{\Theta}&-f\end{pmatrix}=\begin{pmatrix}f&\dfrac{1}{\ell^{\prime}}\bar{\Pi}&0\\ 0&F&\dfrac{1}{\ell^{\prime}}\Pi\\ 0&0&-f\end{pmatrix}+\begin{pmatrix}0&0&0\\ \dfrac{1}{\ell}\Theta&0&0\\ 0&\dfrac{1}{\ell}\bar{\Theta}&0\end{pmatrix}$ (5)
$=\begin{pmatrix}d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta&\dfrac{1}{\ell^{\prime}}(d\bar{\alpha}+\lambda\wedge\bar{\alpha}+\bar{\alpha}\wedge A)&0\\ \dfrac{1}{\ell}(d\beta+\lambda\wedge\beta+A\wedge\beta)&dA+\dfrac{1}{2}[A,A]+\dfrac{1}{\ell\ell^{\prime}}(\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta})&\dfrac{1}{\ell^{\prime}}(d\alpha+A\wedge\alpha+\lambda\wedge\alpha)\\ 0&\dfrac{1}{\ell}(d\bar{\beta}+\bar{\beta}\wedge A+\lambda\wedge\bar{\beta})&-d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\beta}\wedge\alpha\end{pmatrix}$ (6)
with
$F=R+\dfrac{1}{\ell\ell^{\prime}}\phi\quad;\quad$ $\phi=\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta}\quad;\quad f=d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta$
$\Pi=d\alpha+A\wedge\alpha+\lambda\wedge\alpha\quad$ $\text{and}\quad\Theta=d\beta+\lambda\wedge\beta+A\wedge\beta=T+\lambda\wedge\beta\ .$
The Bianchi identities are given by $D\bar{\Omega}=d\bar{\Omega}+\varpi\wedge\bar{\Omega}-\bar{\Omega}\wedge\varpi=0$ which can be split into the identities:
$(D\bar{\Omega})_{-1}=\dfrac{1}{\ell}(d\Theta-F\wedge\beta+\beta\wedge f+A\wedge\Theta-\Theta\wedge\lambda)=0$
$(D\bar{\Omega})_{\mathfrak{so}(3,1)}=\underbrace{dR+[A,R]}_{=0}+\dfrac{1}{\ell\ell^{\prime}}(d\phi+[A,\phi]+\beta\wedge\bar{\Pi}-\Pi\wedge\bar{\beta}+\alpha\wedge\bar{\Theta}-\Theta\wedge\bar{\alpha})=0$
$(D\bar{\Omega})_{\mathbb{R}}=-\mathds{1}\bigl{(}df+\dfrac{1}{\ell\ell^{\prime}}(\bar{\Theta}\wedge\alpha-\bar{\beta}\wedge\Pi)\bigr{)}=0$
$(D\bar{\Omega})_{1}=\dfrac{1}{\ell^{\prime}}(d\bar{\Pi}+\bar{\alpha}\wedge F-f\wedge\bar{\alpha}-\bar{\Pi}\wedge A+\lambda\wedge\bar{\Pi})=0.$
@@
If we impose a normal Cartan geometry such that $\bar{\Omega}_{\mathfrak{p}}=0$ and also consider a gauge where $\lambda=0$, then, as described in [14], $\bar{\Pi}$ and $\bar{\alpha}$ become respectively the Cotton and Schouten tensor while $F$ and $\Theta$ would correspond to the Weyl and torsion tensor. These tensors constitute the building blocks of conformal gravity.
@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
The trace $\operatorname{Tr}$ used to compute the Pontrjagin number associated to $\bar{\Omega}$ is given by
$\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{so}}^{ab}\varphi_{\mathfrak{so}}^{{}^{\prime}cd}+2\varphi_{\mathbb{R}}\varphi^{\prime}_{\mathbb{R}}+2\eta_{ab}(\varphi_{-1}^{a}\varphi_{1}^{{}^{\prime}b}+\varphi_{1}^{a}\varphi_{-1}^{{}^{\prime}b}).$
Taking a linear combination of the Euler and Pontrjagin numbers ($\mathdutchcal{E}$ and $\mathdutchcal{P}$) associated respectively to $F$ and $\bar{\Omega}$ yields the action:
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(F)=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})+yP(F)\Bigr{)}$ (1)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}F^{ab}\wedge F^{cd}-\dfrac{r}{8\pi^{2}}Tr(\bar{\Omega}\wedge\bar{\Omega})-\dfrac{y}{8\pi^{2}}Tr(F\wedge F)\Bigr{)}$
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\underbrace{\phi^{ab}\wedge\phi_{ab}}_{=0}\bigr{)}-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2rf\wedge f\Bigr{)}\biggr{)}$
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{r_{1}s_{1}r_{2}s_{2}}R^{ab}_{\;\;\;r_{1}s_{1}}\beta^{\,c}_{~{}\,\,\mu_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{2}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;r_{2}}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;s_{2}}\eta_{ca}\eta_{db}d^{4}x\bigr{)}$
$\qquad\qquad\quad\quad-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2rf\wedge f\Bigr{)}\biggr{)}.$
By considering the very particular case where [^8] This constraint does not seem to come from a mutation process as defined in Def.2.1. $\alpha=\dfrac{k}{2}\beta$, we get $\Pi=\dfrac{k}{2}\Theta=\dfrac{k}{2}(T+\lambda\wedge\beta)$, $\phi=\ell^{2}\xi$, $F=R+k\xi$ and one arrives at the following action:
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(F)=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})+yP(F)\Bigr{)}=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}(\dfrac{F}{2\pi})+rP(\bar{\Omega})+yP(F)\Bigr{)}$ (2)
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\underbrace{\xi^{ab}\wedge\xi_{ab}}_{=0}\bigr{)}-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}\Theta^{a}\wedge\Theta^{b}-2rf^{2}\Bigr{)}\biggr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}\bigr{)}-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2rd\lambda\wedge d\lambda\Bigr{)}\biggr{)}$
@@
$\qquad\qquad\quad\quad-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2rd\lambda\wedge d\lambda\Bigr{)}\biggr{)}$
and the resulting action for $k=0$ is :
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)R^{ab}\wedge R_{ab}-2rd\lambda\wedge d\lambda\bigr{)}\Bigr{)}.$ (3)
This corresponds to the action (5) for $G=ISO(3,1)\Leftrightarrow k=0$, up to the $\lambda$ kinetic term. However, ignoring the boundary terms leaves the equations of motion unchanged if we keep the same matter Lagrangian as in the preceding example 3.
While for $k\neq 0$ a direct comparison between $S^{\prime}$ (4) and $S_{G}$ (2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{1}\\ \dfrac{(r+y)k}{2\pi^{2}\ell\ell^{\prime}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r+y}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{5}\\ \dfrac{ek^{2}}{8(\pi\ell\ell^{\prime})^{2}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{3}=-2(1+\dfrac{y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{2ek}{(r+y)\ell\ell^{\prime}}\alpha_{3}\\ r+y=\dfrac{2\pi^{2}\ell\ell^{\prime}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (4)

### v5

@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{m}}=\varpi_{0}\oplus\overbrace{\varpi_{1}\oplus\varpi_{-1}}^{\varpi_{\mathfrak{m}}}=$ $\begin{pmatrix}\lambda&0&0\\ 0&A&0\\ 0&0&-\lambda\end{pmatrix}+\varpi_{1}^{a}\begin{pmatrix}0&\dfrac{1}{\ell^{\prime}}\bar{e}_{a}&0\\ 0&0&\dfrac{1}{\ell^{\prime}}e_{a}\\ 0&0&0\end{pmatrix}+\varpi_{-1}^{a}\begin{pmatrix}0&0&0\\ \dfrac{1}{\ell}e_{a}&0&0\\ 0&\dfrac{1}{\ell}\bar{e}_{a}&0\end{pmatrix}$
@@
where $\varpi_{\mathfrak{m}}$ is the soldering form, $A=\dfrac{1}{2}A^{ab}J_{ab}=\dfrac{1}{2}A^{ab}_{\mu}dx^{\mu}\otimes J_{ab}$ corresponds to the spin connection and $\beta=\beta^{a}e_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes e_{a}$ is the tetrad used to define the metric of the tetrad formalism $\tilde{g}=\eta(\beta,\beta)$. $\alpha=\alpha^{a}e_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes e_{a}$ can be interpreted as a secondary tetrad. Finally, $\lambda=\lambda_{\mu}dx^{\mu}$ is the $\mathbb{R}$-valued (dilations) part of the connection. Also, $\bar{\alpha}=\alpha^{a}\bar{e}_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$ and $\bar{\beta}=\beta^{a}\bar{e}_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$.
Remark.
The geometry modeled on the present $8$-dimensional quotient $G/H$ describes in general an $8$-dimensional manifold $\widetilde{\mathscr{M}}$ invariant under Lorentz transformations (elements in $SO(3,1)$) and dilations (elements of the multiplicative group $\mathbb{R}^{+}$) with soldering form $\varpi_{\mathfrak{m}}$. In the next section, we will see how we can obtain a description of a $4$-dimensional submanifold $\mathscr{M}$ of $\widetilde{\mathscr{M}}$ by imposing a condition on the soldering form (the $\mathfrak{m}$-valued part of the Cartan connection).
@@
$\bar{\Omega}$ $=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}=\begin{pmatrix}f&\dfrac{1}{\ell^{\prime}}\bar{\Pi}&0\\ \dfrac{1}{\ell}\Theta&F&\dfrac{1}{\ell^{\prime}}\Pi\\ 0&\dfrac{1}{\ell}\bar{\Theta}&-f\end{pmatrix}=\begin{pmatrix}f&0&0\\ 0&F&0\\ 0&0&-f\end{pmatrix}+\begin{pmatrix}0&\dfrac{1}{\ell^{\prime}}\bar{\Pi}&0\\ \dfrac{1}{\ell}\Theta&0&\dfrac{1}{\ell^{\prime}}\Pi\\ 0&\dfrac{1}{\ell}\bar{\Theta}&0\end{pmatrix}$ (7)
$=\begin{pmatrix}d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta&\dfrac{1}{\ell^{\prime}}(d\bar{\alpha}+\lambda\wedge\bar{\alpha}+\bar{\alpha}\wedge A)&0\\ \dfrac{1}{\ell}(d\beta-\lambda\wedge\beta+A\wedge\beta)&dA+\dfrac{1}{2}[A,A]+\dfrac{1}{\ell\ell^{\prime}}(\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta})&\dfrac{1}{\ell^{\prime}}(d\alpha+A\wedge\alpha+\lambda\wedge\alpha)\\ 0&\dfrac{1}{\ell}(d\bar{\beta}+\bar{\beta}\wedge A-\lambda\wedge\bar{\beta})&-d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\beta}\wedge\alpha\end{pmatrix}$
with:
$\mathfrak{h}\text{-part}\begin{cases}\mathfrak{so}\text{-part:}\quad F=R+\dfrac{1}{\ell\ell^{\prime}}\phi\quad\text{with}~{}\phi=\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta}\\ \\[-11.38109pt] \text{dilation:}\quad f=d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta\end{cases}$
$\mathfrak{m}\text{-part}\begin{cases}\Pi=d\alpha+A\wedge\alpha+\lambda\wedge\alpha,\qquad\bar{\Pi}=d\bar{\alpha}+\bar{\alpha}\wedge A-\bar{\alpha}\wedge\lambda\\ \\[-11.38109pt] \Theta=d\beta+A\wedge\beta-\lambda\wedge\beta=T-\lambda\wedge\beta,\qquad\bar{\Theta}=\bar{T}+\bar{\beta}\wedge\lambda\end{cases}$
(see e.g. [26]). For further use, it is worthwhile to notice that the $[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ contribution to $\bar{\Omega}_{\mathfrak{h}}$ is given by
$\tfrac{1}{2}[\varpi_{\mathfrak{m}},\varpi_{\mathfrak{m}}]=\dfrac{1}{\ell\ell^{\prime}}\begin{pmatrix}\bar{\alpha}\wedge\beta&0&0\\ 0&\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta}&0\\ 0&0&\bar{\beta}\wedge\alpha\end{pmatrix}.$ (9)
The Bianchi identities are given by $D\bar{\Omega}=d\bar{\Omega}+[\varpi,\bar{\Omega}]=0$ and, on account of the identification (1), splits into the identities for each sector
$\mathfrak{h}\text{-part}\begin{cases}\mathfrak{so}\text{-part:}\quad dR+[A,R]+\frac{1}{\ell\ell^{\prime}}\big{(}d\phi+[A,\phi]+\beta\wedge\bar{\Pi}-\Pi\wedge\bar{\beta}+\alpha\wedge\bar{\Theta}-\Theta\wedge\bar{\alpha}\big{)}=0\\ \\[-11.38109pt] \text{dilation:}\quad d(\bar{\alpha}\wedge\beta)+\bar{\alpha}\wedge\Theta-\bar{\Pi}\wedge\beta=0\end{cases}$
$\mathfrak{m}\text{-part}\begin{cases}d\Theta+(A-\lambda)\wedge\Theta+(d\lambda-R)\wedge\beta=0\\ \\[-11.38109pt] d\Pi+(A+\lambda)\wedge\Pi-(d\lambda-R)\wedge\alpha=0.\end{cases}$
@@
If we choose a Cartan geometry such that $\Theta=0$, $f=0$ and Ric$(F)=0$, [11, Chapter 7, Def 2.7] then, as described in [27, p.52], $\bar{\Pi}$ and $\bar{\alpha}$ become, respectively, the Cotton and Schouten tensors, while $F$ would correspond to the Weyl tensor. These tensors constitute the building blocks of conformal gravity.
@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
The trace $\operatorname{Tr}$ used to compute the Pontrjagin number associated to the $\mathfrak{g}$-valued curvature $\bar{\Omega}$ is given by
$\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{so}}^{ab}\varphi_{\mathfrak{so}}^{{}^{\prime}cd}+2\varphi_{\mathbb{R}}\varphi^{\prime}_{\mathbb{R}}+\dfrac{2}{\ell\ell^{\prime}}\eta_{ab}(\varphi_{-1}^{a}\varphi_{1}^{{}^{\prime}b}+\varphi_{1}^{a}\varphi_{-1}^{{}^{\prime}b})=K_{\mathfrak{g}}(\varphi,\varphi^{\prime})/4$
in view of the decomposition (3) and (10) below.
Taking a linear combination of the type (8) made of the invariant polynomial in $F$ (the Pfaffian for the $\mathfrak{so}$-part) and the Pontrjagin number $\mathdutchcal{P}(\bar{\Omega})$ yields the action:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}_{i}}\bigl{(}e\operatorname{Pf}(\dfrac{F}{2\pi})+r\overbrace{\det(\mathds{1}+\dfrac{\bar{\Omega}}{2\pi})}^{P(\bar{\Omega})}+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})\bigr{)}$ (1)
where $\mathscr{M}_{i}$ is either the $8$-dimensional manifold $\widetilde{\mathscr{M}}$ defined by the Cartan geometry $G/H$ or a $4$-dimensional submanifold $\mathscr{M}\subset\widetilde{\mathscr{M}}$ defined by the identification $\alpha=\dfrac{k}{2}\beta$ [^17] This ought to resemble a gauge fixing which amounts to freezing the degrees of freedom of the $\mathfrak{g}_{1}$-part to be those of the $\mathfrak{g}_{-1}$-sector. It can be linked with [11, Proof of Proposition 3.1, p.285] where the two tetrads are taken to be linearly dependent each other. To some extent, this selects a mutated Weyl geometry within the mutated Möbius one. One thus gets $\dim\mathscr{M}=11-7=4$. One may also consider manifolds of other dimensions by imposing other constraints that reduce the number of degrees of freedom of the soldering form $\varpi_{\mathfrak{m}}$.
Remark.
The Pfaffian of $F$, that is $\operatorname{Pf}(\dfrac{F}{2\pi})$ is null if $dim(\mathscr{M}_{i})\neq 4$.
For the time being we focus on the case of an integral on $\mathscr{M}_{i}=\mathscr{M}$ corresponding to the action:
$S_{G}[\varpi]$ $=r\mathdutchcal{P}(\overline{\Omega})+\int_{\mathscr{M}}(e\operatorname{Pf}(\dfrac{F}{2\pi})+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi}))$
$=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}(\dfrac{F}{2\pi})-\dfrac{r}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}\wedge\bar{\Omega})-\dfrac{y}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$ (2)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}F^{ab}\wedge F^{cd}-\dfrac{r}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}\wedge\bar{\Omega})-\dfrac{y}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\underbrace{\phi^{ab}\wedge\phi_{ab}}_{=0}\bigr{)}$
$\qquad\quad-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2(r+y)f\wedge f\Bigr{)}\biggr{)}$
@@
$\qquad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{r_{1}s_{1}r_{2}s_{2}}R^{ab}_{\;\;\;r_{1}s_{1}}\beta^{\,c}_{~{}\,\,\mu_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{2}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;r_{2}}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;s_{2}}\eta_{ca}\eta_{db}d^{4}x\bigr{)}$
$\qquad\qquad\quad\quad-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2(r+y)f\wedge f\Bigr{)}\biggr{)}.$
One can observe that unless $\alpha\propto\beta$ we do not retrieve the Palatini term nor the cosmological constant term.
From the very construction of the action involving polynomials of degree 2 in the curvature $\bar{\Omega}$ the gauge action is invariant under the duality symmetry $\bar{\Omega}\leftrightarrow-\bar{\Omega}$  [^18] Considering instead the action on $\mathscr{M}_{i}=\widetilde{\mathscr{M}}$ one finds the duality symmetry $\bar{\Omega}\leftrightarrow\sqrt[4]{1}\bar{\Omega}$. which can be split as:
$\begin{cases}F\leftrightarrow-F\\ f\leftrightarrow-f\\ \{\dfrac{\Pi}{\ell^{\prime}}\leftrightarrow-\dfrac{\Theta}{\ell}\}\cup\left\{\{\dfrac{\Pi}{\ell^{\prime}}\leftrightarrow-\dfrac{\Pi}{\ell^{\prime}}\}\cap\{\dfrac{\Theta}{\ell}\leftrightarrow-\dfrac{\Theta}{\ell}\}\right\}\end{cases}$ (3)
the action being invariant under each of these transformations. Particular transformations satisfying 3 are:
$\begin{cases}R\leftrightarrow-\dfrac{1}{\ell\ell^{\prime}}(\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta})\\ d\lambda\leftrightarrow-\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta\\ \{\dfrac{1}{\ell}\beta\leftrightarrow-\dfrac{1}{\ell^{\prime}}\alpha\}\cup\left\{\{\dfrac{\alpha}{\ell^{\prime}}\leftrightarrow-\dfrac{\alpha}{\ell^{\prime}}\}\cap\{\dfrac{\beta}{\ell}\leftrightarrow-\dfrac{\beta}{\ell}\}\right\}\end{cases}$ (4)
Thus, upon identifying $\alpha=\dfrac{k}{2}\beta$, the curvature entries reduce to
$\Theta=T-\lambda\wedge\beta,\quad\Pi=\dfrac{k}{2}(T+\lambda\wedge\beta),\quad\phi=k\beta\wedge\bar{\beta}=k\ell^{2}\xi,\quad F=R+\dfrac{k\ell}{\ell^{\prime}}\xi\quad\text{and}\quad f=d\lambda$
and the action (3.2) reads:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\underbrace{\xi^{ab}\wedge\xi_{ab}}_{=0}\bigr{)}-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}\Theta^{a}\wedge\Theta^{b}-2(r+y)f^{2}\Bigr{)}\biggr{)}$
@@
$+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}\bigr{)}-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-2(r+y)d\lambda\wedge d\lambda\Bigr{)}\biggr{)}$
@@
$\qquad\qquad\quad\quad-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-2(r+y)d\lambda\wedge d\lambda\Bigr{)}\biggr{)}.$ (5)
The resulting action for $k=0$ reads:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)R^{ab}\wedge R_{ab}-2(r+y)d\lambda\wedge d\lambda\bigr{)}\Bigr{)}.$ (6)
and corresponds to the action (5) for $G=ISO(3,1)\Leftrightarrow k=0$, up to the $\lambda$ kinetic term. However, ignoring the boundary terms leaves the equations of motion unchanged if we keep the same matter action as in the preceding example 3.
While for $k\neq 0$, a direct comparison between $S^{\prime}$ (4) and $S_{G}$ (3.2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{8\pi^{2}\ell\ell^{\prime}}=\alpha_{1}\\ \dfrac{(r+y)k}{4\pi^{2}\ell\ell^{\prime}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r+y}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{5}\\ \dfrac{ek^{2}}{32\pi^{2}(\ell\ell^{\prime})^{2}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{2k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{4\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{2k}{\ell\ell^{\prime}}\alpha_{3}=-(1+\dfrac{y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{ek}{(r+y)\ell\ell^{\prime}}\alpha_{3}\\ r+y=\dfrac{4\pi^{2}\ell\ell^{\prime}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (7)